OpenStudy (anonymous):

Solve the IVP: dy/dt=te^(t-y) y(0)=5

OpenStudy (jamesj):

Looks separable to me.

OpenStudy (anonymous):

How so?

OpenStudy (jamesj):

e^(t-y) = e^t . e^-y

OpenStudy (anonymous):

Oh duh forgot that property :)

OpenStudy (anonymous):

so separating them looks like this: e^y dy=te^t dt

OpenStudy (anonymous):

then integrate both sides and you get e^y +\[c_{1}\]=te^t-e^t +\[c_{2}\]

OpenStudy (jamesj):

Or more simply, \[ e^y = (t-1)e^x + C \] thus y = ...

OpenStudy (anonymous):

combine c's and get e^y+\[c _{3}\]=te^t-e^t

OpenStudy (jamesj):

I mean (t-1)e^t + C

OpenStudy (jamesj):

\[ e^y = (t-1)e^t + C \] thus y = ...

OpenStudy (anonymous):

y=tln(t-1)+\[c _{4}\]

OpenStudy (jamesj):

No, you can't take the constant outside.

OpenStudy (jamesj):

ln(a + b) does not equal ln a + ln b

OpenStudy (jamesj):

hence \[ y = \ln( (t-1)e^t + C) \]

OpenStudy (anonymous):

okay so it is y=ln(e^t(t-1)+\[c _{3}\])

OpenStudy (anonymous):

and plugging in y(0)=5 gives me \[c _{3}\]=e^5 ???

OpenStudy (jamesj):

Indeed

OpenStudy (anonymous):

so the final answer would be: y=ln(e^t(t-1)+e^5)

OpenStudy (jamesj):

You can verify this by making sure it satisfies the original equation.

OpenStudy (anonymous):

opps \[c _{3}\]=e^5 +1

OpenStudy (anonymous):

and it works :)

OpenStudy (jamesj):

good